Displaying report 1-1 of 1.
Reports until 13:04, Tuesday 08 October 2024
Electronics General
Torrey Cullen - posted 13:04, Tuesday 08 October 2024 - last comment - 21:56, Monday 14 October 2024(11894)
Electronics moving out of b102

[Jeff, Torrey]

We have started moving electronics out of B102. We started with the three Moku's not in use by the RBQ experiment. Two have been mounted in the rack as seen in the photo. The 3rd doesn't have a mount yet, Jeff is emailing liquid instruments about this.

Images attached to this report
Comments related to this report
Torrey Cullen - 13:09, Tuesday 08 October 2024 (11895)

I've disassembled and moved all main electronics associated with the filter cavities to B111. This includes the SHG temperature controller, 1811 15V power sources, RF amplifiers for AOMs (two free standing ones and the two in the chasis Daniel made), HV amplifier for piezo control, a spare function generator, and all 7 photo detector power sources. A permanent spot for this has not yet been found.

Jeffrey Wack - 16:07, Friday 11 October 2024 (11901)

Moku is sending us two sets of rack mounting hardware (for free)

Torrey Cullen - 14:21, Monday 14 October 2024 (11902)

[Jeff, Daniel, Torrey]

We have wrapped the optical sleds in Mylar and are ready to move. With Sander and Alex out, and Ian still injured, we need to wait til Daniel Jeff and I are all free at the same time to make the final move. Tentatively, this will be Tuesday. I may ask JC to join as well. Hopefully moving optics soon!

Jeffrey Wack - 21:56, Monday 14 October 2024 (11904)

All 4 Moku Pros now have rack mounting fins installed

Displaying report 1-1 of 1.